# Patch-Consistent Optical Translation Across Sensors: Large-Scale Denoising Diffusion with Heterogeneous Change Detection as a Use Case

### v2

@@ III Dataset
IV Experiments
IV-A Image Quality Metrics and Ablation Experiments
@@
To effectively demonstrate the performance and advantages of the proposed I2I algorithm, a comparison in terms of image quality metrics is proposed. Table I displays the values of three image quality/comparison metrics: Frechét Inception Distance (FID), mean Peak Signal-to-Noise Ratio (PSNR), and mean Learned Perceptual Image Patch Similarity (LPIPS). To obtain these values, we use the test set described in Section III. The FID [Heusel2017] measures the similarity between two data sets of images—usually a data set of generated images and one of reference images—by computing the distance between their feature representations extracted by a pre-trained Inception network. A lower FID indicates better feature alignment of the generated images with the reference data set. The PSNR [Hore2010] measures the reconstruction error between two images in terms of the ratio of the maximum possible pixel value and the mean squared error. A higher PSNR indicates a lower pixel-wise reconstruction error. It is important to note that PSNR is not a reliable metric for perceptual quality, as it does not account for the human visual system and the semantic content of the images [Saharia2021]. Nonetheless, in the presented results, we compute the mean PSNR (mPSNR) of all Planet and Synthetic Planet corresponding patches, to further enrich the discussion. The LPIPS [Zhang2018LPIPS] measures the perceptual similarity between a pair of images by computing the distance between their feature representations extracted by a pre-trained deep neural network. A lower LPIPS indicates higher perceptual similarity and better visual quality. LPIPS is aligned with human perception, as it captures the semantic and structural differences between images [Zhang2018LPIPS]. As with the PSNR, we compute the mean LPIPS (mLPIPS) of all corresponding Planet and Synthetic Planet patches. In a patch-wise comparison, the mLPIPS metric is better suited as a comparison metric than FID since its calculation is done on a patch-by-patch basis instead of focusing on the whole test set at once. Thus, since our tests involve direct comparison between corresponding images, mLPIPS is also employed.
@@ IV-B Visual Examples
In a direct comparison between Sentinel-II, Planet Dove, and the synthetic Planet Dove image generated by the proposed method, Figure 5 presents a region from the city of Tlaquepaque, Mexico. Each of the three rows of figures shows a different view of the city, with different scales, to better compare different aspects of the images. The three columns show images from Planet Dove, Sentinel-II, and images generated by the proposed method, respectively. The first row shows a far view of the city that depicts multiple complex urban features: houses, buildings, storage facilities, factories, roads, and scattered vegetation. When looking at the quality of the figures, Figure 5a clearly shows great contrast between features due to it coming from a higher spatial resolution satellite when compared to Figure 5b. Figure 5c is the synthetic high-resolution image generated with the proposed model using Figure 5b as input. It exhibits a higher contrast between features, compared to the Sentinel-II version, while maintaining feature consistency between patches to the point that it becomes almost impossible to distinguish individual patches in the full images. The same can be said about Figure 5f, which shows a close-up view of a region located around the center of Figure 5c, where homogeneity is still observed, and the super-resolution aspect of the method starts to be noticeable. That is, high-frequency information not clearly observable in Figure 5e gets enhanced by the algorithm, resulting in Figure 5f. Finally, the last row depicts a highly zoomed-up area of the city, where individual residential buildings can be visualized. It is clearly noticeable in Figure 5g how the higher spatial resolution information provided by the Planet Dove sensor enables the visualization of details otherwise imperceptible in the Sentinel-II image (Figure 5h), as well as properly separates features coming from different elements in the image, e.g., roads are easily distinguishable from buildings, whereas in 5h the pixels from narrow roads are interleaved with pixels from buildings. This drawback of Sentinel-II images is alleviated by the proposed method, as can be observed in Figure 5i. There, pixel-level details are almost as clear and as distinguishable as in Figure 5g. In addition to this, the synthetic image does not show visible undesirable hallucinations, which enables change detection tasks.
@@
Figures 6 and 7 respectively show images from the selected regions of Beirut and Austin. Their first, second, third and fourth rows individually show pre-event Sentinel-II, pre-event Synthetic Planet Dove (Regression + $\mathbf{W}$), pre-event Synthetic Planet Dove (Proposed), pre-event Planet Dove, post-event Planet Dove, and the change detection ground truth obtained by applying Algorithm 5, which is discussed in Section V, using Planet Dove pre and post event images. Instead of manual annotations, we use this ground truth since our I2I translation method aims to closely match low-resolution input to the style and resolution of high-resolution images, specifically Planet Dove in our experiments. In both regions, the tonality of the synthetic images generated using the proposed method (Figs. 6c and 7c) closely resemble those of the Planet pre-event images (Figs. 6d and 7d). Also, when they are compared with the pre-event Sentinel-II images used as inputs to the DDM-based method (Figs. 6a and 7a), it is easily observable how they have an improved apparent resolution and contrast. It is also noticeable how no perceptible undesired hallucinated features are present, favoring their use for change detection. Compared to the Regression + $\mathbf{W}$ synthetic images (Figs. 6b and 7b), our method delivered images with higher contrast and visual resolution, which translated into images that better match the characteristics of Planet Dove imagery.
@@
V Change Detection as a Use Case
@@
V-A Rationale for Selective Comparisons
@@
It is important to note that the tested change detection algorithm does not necessarily extract the full potential of our image translation method, as it contains steps that do not fully utilize the added visual resolution to our advantage. Additionally, using this specific simple HCD method with images generated by the proposed method does not necessarily produce better HCD results when compared to other methods such as ShuffleMixer, SRDenseNet, and SwinIR. These methods have performed poorly in super resolution but very well in patch-wise feature consistency, and Algorithm 5 does not make use of the higher resolution features produced by our image translation method. However, employing a more advanced change detection algorithm is beyond the scope of this paper. The HCD experiments are intended to demonstrate that our proposed method could be potentially useful for change detection. A more advanced learning-based change detection procedure using the produced synthetic images will be explored in future work. Furthermore, Regression + + $\mathbf{W,C}$ is trained with the same neural network backbone as the proposed method. By doing so, we effectively showcase the advantages of our diffusion-based method under consistent conditions, ensuring a clear and concise presentation of our results.
V-B Procedure
@@
V-C Experiments
Two pairs of Sentinel-II pre-event and Planet Dove post-event images from two different cities were selected for a change detection comparison experiment. We also use their respective pre-event Planet Dove images to produce a change detection ground truth. The Sentinel-II pre-event images are part of the test set mentioned in Section III. The first pair of images is in Beirut, Lebanon, surrounding the port region. It has been extracted from the northern part of the images in Figure 1, where no clouds are present. This image has been considered because it showcases multiple different types of changes, making targetless change detection useful. Moreover, it is a good example to test the robustness of the proposed method. The second pair of images is in Austin, USA, where many construction and terrain-related changes are present. The metrics used henceforth are the detection rate (DR) (recall) and the false alarm rate (FAR), which is defined as the ratio between the number of pixels misclassified as changes and the total number of unchanged pixels. These are metrics that properly quantify the trade-off between over and under detection.
V-C1 Visual Evaluation
@@ V-C2 Quantitative Evaluation
VI CONCLUSION
@@
Appendix A Related Works
A-1 Denoising Diffusion Models
A-2 Image-to-Image Translation
Despite the advancements in image-to-image translation within the context of remote sensing, existing literature predominantly focuses on the generation of additional training data for direct use in posterior deep learning algorithms [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], or in SAR to Optical/Optical to SAR translation [CDLi2021, I2IZhang2023, I2ISeo2023]. These methods, while highly valuable, do not address the challenge of translating large-scale images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images. This gap is significant as it pertains to the fidelity of the translated images, where the preservation of salient features without introducing non-existent elements is crucial. Our work pioneers in this direction by proposing a novel framework that not only translates images from varying-resolution optical sensors but also guarantees the consistency of features. This approach mitigates the common issue of model hallucinations, where the generation of artifacts or the alteration of key elements could lead to misinterpretations, especially in critical applications like environmental monitoring and urban planning. By maintaining strict adherence to feature consistency, our method ensures that the output images faithfully represent the true characteristics of the input.
Appendix B Implementation Details
@@
We use the Huber loss function [meyer2020alternative] with its parameter $\delta$ set to $0.5$ since it has been shown to reduce the impact of outliers in the gradients while maintaining the benefits of the mean square distance for non-outliers. Implementation details for the Huber loss are presented in Appendix C-A3. RAdam optimizer [liu2021variance] has been chosen due to its proven training speed and its low sensitivity to the learning rate choice. The training ran for $31$ epochs, which was more than enough for the loss function to converge completely.
@@
For the performed experiments, multiple hyperparameters not directly related to the model setup have been chosen. They are presented in Table III, along with their purpose. Some choices were based on a grid search, and others were based on the articles that defined them in the first place. Moreover, two recently proposed training techniques for denoising diffusion models were included in our experiments: Consistency Diffusion Loss and Min-SNR-$\gamma$ Weighting. They are further discussed in Appendices C-D and C-E. The training and experiments were performed using one NVIDIA RTX A6000 GPU.
B-A Model Pruning
@@
B-B Discussion on the Selection of Hyperparameters
@@
The training process itself was meticulously monitored, with the batch size and learning rate being set to values that allowed for stable and complete convergence of the loss function within 31 epochs. The use of the Huber loss function and the RAdam optimizer further contributed to the robustness and speed of the training, as detailed in Appendix C-A3.
@@
Appendix C Fundamental Concepts
C-A Conditional Denoising Diffusion Probabilistic Models
C-A1 Forward Diffusion
@@ C-A3 Training Objective
C-B Denoising Diffusion Implicit Models
@@
C-C Classifier-Free Guidance
@@
C-D Consistent Diffusion Models

### v3

@@ IV Dataset
V Experiments
V-A Image Quality Metrics and Ablation Experiments
@@
To effectively demonstrate the performance and advantages of the proposed I2I algorithm, a comparison in terms of image quality metrics is proposed. Table I displays the values of three image quality/comparison metrics: Frechét Inception Distance (FID), mean Peak Signal-to-Noise Ratio (PSNR), and mean Learned Perceptual Image Patch Similarity (LPIPS). To obtain these values, we use the test set described in Section IV. The FID [Heusel2017] measures the similarity between two data sets of images—usually a data set of generated images and one of reference images—by computing the distance between their feature representations extracted by a pre-trained Inception network. A lower FID indicates better feature alignment of the generated images with the reference data set. The PSNR [Hore2010] measures the reconstruction error between two images in terms of the ratio of the maximum possible pixel value and the mean squared error. A higher PSNR indicates a lower pixel-wise reconstruction error. It is important to note that PSNR is not a reliable metric for perceptual quality, as it does not account for the human visual system and the semantic content of the images [Saharia2021]. Nonetheless, in the presented results, we compute the mean PSNR (mPSNR) of all Planet and Synthetic Planet corresponding patches, to further enrich the discussion. The LPIPS [Zhang2018LPIPS] measures the perceptual similarity between a pair of images by computing the distance between their feature representations extracted by a pre-trained deep neural network. A lower LPIPS indicates higher perceptual similarity and better visual quality. LPIPS is aligned with human perception, as it captures the semantic and structural differences between images [Zhang2018LPIPS]. As with the PSNR, we compute the mean LPIPS (mLPIPS) of all corresponding Planet and Synthetic Planet patches. In a patch-wise comparison, the mLPIPS metric is better suited as a comparison metric than FID since its calculation is done on a patch-by-patch basis instead of focusing on the whole test set at once. Thus, since our tests involve direct comparison between corresponding images, mLPIPS is also employed.
@@ V-B Visual Examples
In a direct comparison between Sentinel-II, Planet Dove, and the synthetic Planet Dove image generated by the proposed method, Figure 5 presents a region from the city of Tlaquepaque, Mexico. Each of the three rows of figures shows a different view of the city, with different scales, to better compare different aspects of the images. The three columns show images from Planet Dove, Sentinel-II, and images generated by the proposed method, respectively. The first row shows a far view of the city that depicts multiple complex urban features: houses, buildings, storage facilities, factories, roads, and scattered vegetation. When looking at the quality of the figures, Figure 5b clearly shows great contrast between features due to it coming from a higher spatial resolution satellite when compared to Figure 5a. Figure 5c is the synthetic high-resolution image generated with the proposed model using Figure 5a as input. It exhibits a higher contrast between features, compared to the Sentinel-II version, while maintaining feature consistency between patches to the point that it becomes almost impossible to distinguish individual patches in the full images. The same can be said about Figure 5f, which shows a close-up view of a region located around the center of Figure 5c, where homogeneity is still observed, and the super-resolution aspect of the method starts to be noticeable. That is, high-frequency information not clearly observable in Figure 5d gets enhanced by the algorithm, resulting in Figure 5f. Finally, the last row depicts a highly zoomed-up area of the city, where individual residential buildings can be visualized. It is clearly noticeable in Figure 5h how the higher spatial resolution information provided by the Planet Dove sensor enables the visualization of details otherwise imperceptible in the Sentinel-II image (Figure 5g), as well as properly separates features coming from different elements in the image, e.g., roads are easily distinguishable from buildings, whereas in 5g the pixels from narrow roads are interleaved with pixels from buildings. This drawback of Sentinel-II images is alleviated by the proposed method, as can be observed in Figure 5i. There, pixel-level details are almost as clear and as distinguishable as in Figure 5h. In addition to this, the synthetic image does not show visible undesirable hallucinations, which enables change detection tasks.
@@
Figures 6 and 7 respectively show images from the selected regions of Beirut and Austin. Their first, second, third, and fourth rows individually show pre-event Sentinel-II, pre-event Synthetic Planet Dove (Regression + $\mathbf{W}$), pre-event Synthetic Planet Dove (Proposed), pre-event Planet Dove, post-event Planet Dove, and the change detection ground truth obtained by applying Algorithm 5, which is discussed in Section V-C, using Planet Dove pre and post-event images. Instead of manual annotations, we use this ground truth since our I2I translation method aims to closely match low-resolution input to the style and resolution of high-resolution images, specifically Planet Dove in our experiments.
In both regions, the tonality of the synthetic images generated using the proposed method (Figs. 6c and 7c) closely resembles those of the Planet pre-event images (Figs. 6d and 7d). Also, when compared with the pre-event Sentinel-II images used as inputs to the DDM-based method (Figs. 6a and 7a), it is easily observable how they have an improved apparent resolution and contrast. It is also noticeable how no perceptible undesired hallucinated features are present, favoring their use for change detection. Compared to the Regression + $\mathbf{W}$ synthetic images (Figs. 6b and 7b), our method delivered images with higher contrast and visual resolution, which translated into images that better match the characteristics of Planet Dove imagery.
The few patching brightness mismatches can be easily alleviated by performing inference with overlapping patches, as done in [wan_psc_2024]. However, this would greatly increase the numerical complexity of the inference, e.g. by a factor of 4 if the overlap is 1/2 of the size of the patch, or by a factor of 16 if the overlap is 3/4 of the size of the patch. The reverse diffusion process is costly because it needs numerous model forward passes to infer a single patch. Therefore, we do not explore this possibility in our paper.
@@
V-C Change Detection as a Use Case
@@
V-C1 Rationale for Selective Comparisons
@@
It is important to note that the tested change detection algorithm does not necessarily extract the full potential of our image translation method, as it contains steps that do not fully utilize the added visual resolution to our advantage. Additionally, using this specific simple HCD method with images generated by the proposed method does not necessarily produce better HCD results when compared to other methods such as ShuffleMixer, SRDenseNet, and SwinIR. These methods have performed poorly in super resolution but very well in patch-wise feature consistency, and Algorithm 5 does not make use of the higher resolution features produced by our image translation method. However, employing a more advanced change detection algorithm is beyond the scope of this paper. The HCD experiments are intended to demonstrate that our proposed method could be potentially useful for change detection. A more advanced learning-based change detection procedure using the produced synthetic images will be explored in future work. Furthermore, Regression + $\mathbf{W,C}$ is trained with the same neural network backbone as the proposed method. By doing so, we effectively showcase the advantages of our diffusion-based method under consistent conditions, ensuring a clear and concise presentation of our results.
V-C2 Procedure
@@
V-C3 HCD Experiments
Two pairs of Sentinel-II pre-event and Planet Dove post-event images from two different cities were selected for a change detection comparison experiment. We also use their respective pre-event Planet Dove images to produce a change detection ground truth. The Sentinel-II pre-event images are part of the test set mentioned in Section IV. The first pair of images is in Beirut, Lebanon, surrounding the port region. It has been extracted from the northern part of the images in Figure 1, where no clouds are present. This image has been considered because it showcases multiple different types of changes, making targetless change detection useful. Moreover, it is a good example to test the robustness of the proposed method. The second pair of images is in Austin, USA, where many construction and terrain-related changes are present. The metrics used henceforth are the detection rate (DR) (recall) and the false alarm rate (FAR), which is defined as the ratio between the number of pixels misclassified as changes and the total number of unchanged pixels. These are metrics that properly quantify the trade-off between over and under detection.
Visual Evaluation
@@ V-C3 HCD Experiments
VI Discussion
The results of our study demonstrate the efficacy of leveraging Denoising Diffusion Implicit Models (DDIM) for large-scale optical image translation. Our approach successfully addresses the limitations of existing Image-to-Image (I2I) methods, particularly in handling large-scale multi-patch imagery. By focusing on super-resolving low spatial resolution images into high-resolution equivalents, our method achieves significant improvements in both radiometric accuracy and feature representation.
One of the key findings is the ability of our method to maintain uniformity across hundreds of patches, which is crucial for applications such as heterogeneous change detection (HCD). The comparative study against the standard classifier-free guided DDIM framework and other leading methods highlights the superior performance of our approach in terms of domain adaptation and artifact reduction. The HCD tasks in urban settings like Beirut, Lebanon, and Austin, USA, further validate the practical applicability of our method.
However, it is important to highlight that the simple change detection algorithm employed in our experiments does not fully exploit the enhanced visual resolution provided by our image translation method. This limitation is partly due to the algorithm’s design, which does not leverage the higher-resolution features effectively. Consequently, while our method shows promise, it does not necessarily outperform other methods, such as ShuffleMixer, SRDenseNet, and SwinIR, in HCD tasks using the chosen change detection algorithm. Despite their poor super-resolution performance, these methods maintain patch-wise feature consistency. The HCD experiments in this study are intended to demonstrate the potential utility of our approach rather than to establish it as the definitive solution for change detection. Future work will explore more advanced learning-based change detection procedures that can fully utilize the synthetic images generated by our method.
While our method effectively handles large-scale imagery, the computational complexity remains challenging. Future work will also explore optimization techniques to reduce the computational load without compromising the model’s performance. Additionally, extending our approach to other types of remote sensing data, such as radar or hyperspectral images, could broaden its applicability and impact. Another potential direction for future work is integrating our method with real-time processing systems. This would enable the rapid analysis of remote sensing data, particularly valuable in time-sensitive applications such as disaster response and environmental monitoring. Furthermore, incorporating advanced machine learning techniques, such as reinforcement learning, could enhance the adaptability and robustness of our method in diverse and dynamic environments.
In conclusion, this paper presented a novel and effective DDM-based approach for large-scale optical image translation addressing key challenges. The results demonstrate significant improvements in image quality and consistency, paving the way for more accurate and reliable remote sensing applications.
VII Conclusion
@@
Appendix A Implementation Details
@@
We use the Huber loss function [meyer2020alternative] with its parameter $\delta$ set to $0.5$ since it has been shown to reduce the impact of outliers in the gradients while maintaining the benefits of the mean square distance for non-outliers. Implementation details for the Huber loss are presented in Appendix B-A3. RAdam optimizer [liu2021variance] has been chosen due to its proven training speed and its low sensitivity to the learning rate choice. The training ran for $31$ epochs, which was more than enough for the loss function to converge completely.
@@
For the performed experiments, multiple hyperparameters not directly related to the model setup have been chosen. They are presented in Table III, along with their purpose. Some choices were based on a grid search, and others were based on the articles that defined them in the first place. Moreover, two recently proposed training techniques for denoising diffusion models were included in our experiments: Consistency Diffusion Loss and Min-SNR-$\gamma$ Weighting. They are further discussed in Appendices B-D and B-E. The training and experiments were performed using one NVIDIA RTX A6000 GPU.
A-A Model Pruning
@@
A-B Discussion on the Selection of Hyperparameters
@@
The training process itself was meticulously monitored, with the batch size and learning rate being set to values that allowed for stable and complete convergence of the loss function within 31 epochs. The use of the Huber loss function and the RAdam optimizer further contributed to the robustness and speed of the training, as detailed in Appendix B-A3.
@@
Appendix B Fundamental Concepts
B-A Conditional Denoising Diffusion Probabilistic Models
B-A1 Forward Diffusion
@@ B-A3 Training Objective
B-B Denoising Diffusion Implicit Models
@@
B-C Classifier-Free Guidance
@@
B-D Consistent Diffusion Models